OpenStudy (anonymous):

what is 1 and two-fifths times 4 and two-thirds equals

OpenStudy (anonymous):

plz someone help

OpenStudy (shamil98):

\[\huge 1 \frac{ 2 }{ 5 } \times 4\frac{ 2 }{ 3 }\] convert them to improper fractions first. \[\huge \frac{ 7 }{ 5 } \times \frac{ 14 }{ 3 }\] then multiply the terms \[\huge \frac{ 7 \times 14 }{ 5 \times 3 }\]

OpenStudy (anonymous):

and the answer is dundundun

OpenStudy (anonymous):

drum role plz

OpenStudy (shamil98):

Do the multiplication.

OpenStudy (anonymous):

ugh to hard

OpenStudy (shamil98):

what is 7 x 14?

OpenStudy (anonymous):

hmm give me a minute to thinkl

OpenStudy (anonymous):

98

OpenStudy (anonymous):

is that corect

OpenStudy (anonymous):

:)

OpenStudy (praxer):

98/15